# Curiosity Driven Exploration to Optimize Structure-Property Learning in Microscopy

### v1

@@ II Introduction
Here, we present alternate methods relying on surrogate models of error prediction, which we term curiosity-driven exploration, analogous to the usage of the term in reinforcement learning zheng2021episodic ; thiede2022curiosity ; pathak2017curiosity . These methods are based on standard deep neural networks with an encoder-decoder structure that have been employed in the past to predict spectra from images (Im2spec) and images from spectra (Spec2im)kalinin2021toward . When the goal is to minimize the loss of an Im2spec or Spec2im model, the optimal scalarizer function is difficult, if not impossible to find. As a solution, we instead determine which spectra to measure by training an auxiliary network to predict Im2spec reconstruction error. The curiosity-driven approach involves sampling regions with high values of the predicted error, so as to attempt to rapidly reduce the error of these models.
The paper presents two workflows: The first consists of an ensemble of Im2spec models that is used for spectral prediction, followed by an error model that trains on the spectral mismatch error. In the second method, the error model utilizes the latent space embeddings of an autoencoder to correlate with spectral mismatch. These algorithms, inspired by curiosity-driven reinforcement learning, actively sample spectra for which the structure-property relations have not yet been learned. We first demonstrate and optimize the efficacy of our methods on a pre-acquired dataset. Finally, we implement an algorithm on an atomic force microscope (AFM) to actively learn structure-property relationships in a ferroelectric thin film and discuss possible extensions.
Figure 1: Description of the spectral search method on pre-acquired dataset. (a) Spatial dependence of the spectral property that is correlated with the sample region shown on the left. The patches on the sample image serve as the structural inputs that is correlated to the spectral output. (b) An ensemble of different Im2spec models are trained on an initial set of data. (c) The best model is then used to predict the spectral output corresponding to the training image inputs. (d) Error model consists of the encoder and the latent embedding part of the Im2spec model. This is conjoined with a decoder that is used to train with the spectral mismatch L1-error. The parameters of the encoder (and the latent embedding) are frozen while training the error model. (e) The error model is used to prediction for the image patches across the sample region. The next points are decided using the acquisition function, and incorporated into the training set for the subsequent iteration.
@@ III Results and Discussions
The structure-property relations described in this work correlate to the ferroelectric response of PbTiO3 samples measured using band excitation piezoresponse microscopy (BE-PFM). We use the structure information from the piezoresponse imaging data acquired from the AFM scanned images, while the property is measured using the spectra collected using band excitation piezoresponse spectroscopy (BEPS) data. This dataset is very similar to one captured and published earlier, and details about the measurement can be found elsewhereslautin2024bayesian .
Fig 1 illustrates the active learning workflow described in this section. Fig 1(a) shows sample dataset which shows spatial dependence of the local structure and its influence on the observed spectrum. Here, the local structure, indicated by the square patch, influences the spectrum measured in that region. We initially start by considering a training set where the inputs are the image patches (each patch of size (16 x 16) pixels) while the outputs are the spectra (256 points) corresponding to each patch. This initial set is used to train Im2spec models.
In this workflow, we use an ensemble of Im2spec models to offer flexibility for variations in the training data, as shown in Fig 1(b). While the models are primarily based on the convolutional networks, variations in the architecture and the hyperparameters have been introduced to enable wider adaptability. A brief description of the encoder architectures used in the model set is provided in Table 1. We designate the size of the latent dimension as three. An initial dataset of the image patches is used to train the Im2spec models. Once trained, the best model of the ensemble is chosen based on the minimum validation loss.
@@
The selected model is then used to predict the spectral output on the image inputs that were previously used for training (as shown in Fig 1(c)). This prediction is compared with the original spectrum, and the mismatch error is assigned to every image within the training set. We use the L1 loss to quantify the spectral mismatch in this method. Fig 1(d) shows the error model that consists of the Im2spec-encoder (which includes the latent embedding layer) conjoined with a different set of decoder layers. During the error model training, the encoder part of the model is frozen while the decoder weights are updated.  The next step involves the error prediction for the entire set of image patches across the sample region, as shown in Fig 1(e). The error predictions are used to compute the acquisition function to determine and sample the next set of spectral points in an iterative active learning fashion.
Our studies show that the best Im2spec model does not change frequently with minor changes in the training data set. Our code allows for the assignment of a probability associated with ensemble training to minimize unnecessary training procedures. In the results described in this section, we perform ensemble training randomly over 20% of the iterations (and the starting iteration). The remaining iterations involve model training using the pre-determined best-Im2spec-model.
The acquisition function used in this method is an empirical equation and is given as:
@@
where $L_{j}$ is the L1 error normalized in the range [0, 1]. The $\beta$ parameter controls the rate of exploration and exploitation, while the prefactor $\lambda$ controls the smoothness of the acquisition function (higher $\lambda$ indicates better smoothness). The acquisition function varies monotonically with the error values for $\beta$ = 1 and inversely for $\beta$=0. This allows us to tune sampling from exploitation to exploration as we increase the beta hyperparameter from 0 to 1. In the results described in this section, we study model exploration for different $\beta$ values while maintaining $\lambda$ = 0.1.
@@
The workflow starts with an initial dataset consisting of 245 image-spectrum pairs (20% of the total dataset). Each iteration consists of two model training events - the im2spec ensemble models and the error model. As shown in Fig 1(e), we obtain the prediction of the error values at the end of each iteration. We use the acquisition to sample the next batch of ten points in every iteration. In the results described in this section, we study the model behavior over twenty iterations totaling two hundred explored acquisitions.
Fig 2 shows the workflow results where we test the model for varying values of the $\beta$ parameter in the range [0, 1]. The figures in panel Fig 2(a)-(c) show the the results for $\beta$ = 0, which samples the points associated with low-error values. Fig 2(a) shows the posterior error prediction at the end of iterations. The red scatter points indicate the points of low error values selected by the acquisition function for subsequent sampling. Fig 2(b) shows the points that indicate all acquisitions at the end of the active learning process overlaid on the topographic image. Fig 2(c) shows the statistics of the determined L1 error across all iterations. Given the conservative nature of the exploration associated with $\beta$ = 0, the cumulative error of the points selected by the active learning process is lower than randomly sampled images.
Next, we study the effect of dominant exploration with $\beta$= 1, and the results are shown in Figs. 2(d)-(f). The higher value of the $\beta$ parameter seeks out sample regions with the highest fitting error. This is evident in Fig 2(d), which shows the points selected for sampling in areas with high predicted error. In contrast to the points acquired in Fig 2(b) (for $\beta$ =0), Fig 2(e) shows areas where the points are sampled at complementary locations across the sample region. A closer observation reveals higher sampling in the interfacial regions around the domain wall, where the structure-property correlations are not straightforward. The curious, explorative nature of the model is reflected in Fig 2(f), which shows higher errors associated with the sampled locations.
Figure 2: Results of the error prediction model in active learning on the pre-acquired BEPS data. Panel (a)-(c) shows the results of the model for beta = 0. (a) Error prediction across the sample region. scattered points indicate points for subsequent acquisitions. All scale bars indicate a length of 100 nm. (b) shows the distribution of acquisitions over 20 iterations (200 points) overlayed on topographic image of the sample. (c) shows the distribution of the L1-error for the acquired spectra. Panel (d)-(f) shows the similar results for active learning-based sampling with $\beta$ = 1, while figures (g)-(i) show the results for active learning based sampling for $\beta$ = 0.5.
@@
While we see extreme examples of exploitation and exploration for $\beta$ = 0 and 1, respectively, intermediate values of the $\beta$ can be used to balance exploration and exploitation. Fig (g)-(i) shows the results for $\beta$ = 0.5. This samples regions that correspond to both higher and lower values of the error prediction. This is reflected in the intermediate range of the mean error shown in Fig 2(i).
@@
In an encoder-decoder model, the latent representations determine the efficiency of the reconstruction. We study the latent embeddings that bridge the two different models to gain insights into the workings of the error model and to interpret the essential features that determine the model output. The latent distributions of the model predictions are described in Fig 3 for the active learning process across different values of $\beta$. Fig 3(a)-(c) represents the latent space distributions for $\beta$ = 0. The conservative nature associated with it is seen as the exploration points are localized at the high-density region of the latent space. Further, to correlate the latent encodings to points in the real space, we use a k-means cluster on the latent space, as shown in Fig 3(b), and these are mapped to the real-space coordinates shown in Fig 3(c). A similar analysis is performed for exploration related to $\beta$ = 1, shown in Fig 3(d)-(f). Here, higher exploration has resulted in a dispersed latent distribution. Further exploration points are comparatively sampled in the sparse region of the latent space. In the real space mapping (Fig 3(f)), this translates to acquisition in the complementary areas and corroborates with the data shown in Fig 2(e). However, for intermediate values of the $\beta$ = 0.5, we see sampling across the distribution of the latent space, indicating a balance of exploration and exploitation.
@@
Figure 3: Latent space representations for the active learning-based sampling. Panel (a)-(c) shows the results of the model for beta = 0. (a) Latent distribution of the model prediction for all image-patch inputs. (b) depicts the latent distribution clustered into three components using k-means clustering. (c) Correlation to the real-space coordinates corresponding each of the points in the latent space. Panel (d)-(f) shows the similar results for active learning-based sampling with $\beta$ = 1 while figures (g)-(i) show latent distribution results for active learning-based sampling for $\beta$ = 0.5
@@
These results describe the error prediction methods in conjunction with the acquisition function where the $\beta$ parameter is used to control exploration. The methodology is denoted as a curiosity-driven for $\beta$ = 1 where the model actively searches for the regions of higher predicted error to better learn diverse structure-spectral correlations.
@@
Further, the latent space analysis emphasizes the importance of the latent distributions in its ability to reconstruct the spectra and predict the reconstruction error. These embeddings serve as compressed, structured representations of input data, capturing essential features of the input images. Given this knowledge, in the subsequent section, we implement a generalized methodology to extract latent representations from an autoencoder while efficiently sampling points from the latent space for active learning based acquisitions.
@@ III.2 Autoencoder-based error model
Figure 4: Diagram of Curiosity Algorithm implementation with Im2spec
This section describes the autoencoder based error model. Here we use a similar experimental dataset - PFM-based experimental data on a 200nm (110) $PbTiO_{3}$ thin film sample grown on $SrTiO_{3}$. The structural information contained in the image patches (patch size (11$\times$11 pixels)) with switching spectroscopy spectra (spectrum length:64) captured at low frequency (off-resonance) using an interferometric atomic force microscope from Oxford Instruments (Vero). The baseline curiosity algorithm works as follows: After the sample is imaged an autoencoder is trained on all image patches. Then spectra are acquired on a small number of random initialization points, which are used to train the Im2spec model. The error predictor is then trained on the image patch latent encodings and the Im2spec mean squared error (MSE) for the initial points. Then, a forward pass through the error predictor is performed for all image patch encodings. The spectra of the point with the highest predicted error is then sampled. This continues iteratively where Im2spec and the error predictor are trained on updated training dataset. The overall workflow is illustrated in Fig 4. It should be noted that the error predictor model utilizes dropout to provide an estimate of the uncertainty on the prediction.
@@
Figure 6: Minimum loss achieved by Im2spec with Curiosity Algorithm vs Random sampling.
Figure 7: MSE for ten highest error points achieved by Im2spec with MC Dropout Curiosity Algorithm vs Random sampling.
Figure 8: Trial of Curiosity Algorithm real-time on PFM microscope for Spec2im and Im2spec. (a) Exploration path of Spec2im Curiosity Algorithm, (b) predicted Spec2im error after final measurement iteration, (c) a polarization ground-truth image-patch, and (d) corresponding Spec2im prediction. (e) Exploration path of Im2spec Curiosity Algorithm, (f) predicted Im2spec error after final measurment iteration, (g) a ground-truth hysteresis loop and corresponding Im2spec prediction. Scale bar in the images indicate a length of 100 nm.
@@
In addition, we tested a modified curiosity algorithm which, in addition to latent space exploration reward, samples based on Im2spec Monte-Carlo Dropout (MCD) uncertainty during the exploration phase. While the addition of MCD uncertainty did not directly improve Im2spec loss, it reduced the Im2spec MSE for the ten highest error points (Fig. 7). This behavior suggests that enhancing exploration with MCD helps train Im2spec on points with poorly understood structure property relationships, but are not abundantly represented in the sample data, as opposed to points with low error, but are highly represented in the sample data. It should be noted that one of the challenges of this algorithm is that there may exist points that continue to contain high errors regardless of the number of training data points, if there are minimal structure-property correlations in these points (for example, if there is only noise in these areas). For such instances, the algorithm should be modified to avoid trapping in these learning plateaus, ans strategies can include either direct human intervention, injected noise in the action space, or simple methods such as avoiding similar image patches to past samples if the loss is not decreasing beyond a simple threshold.

### v2

@@ II Introduction
Here, we present alternate methods relying on surrogate models of error prediction, which we term curiosity-driven exploration, analogous to the usage of the term in reinforcement learning zheng2021episodic ; thiede2022curiosity ; pathak2017curiosity ; burda2018exploration . These methods are based on standard deep neural networks with an encoder-decoder structure that have been employed in the past to predict spectra from images (Im2spec) and images from spectra (Spec2im)kalinin2021toward . When the goal is to minimize the loss of an Im2spec or Spec2im model, the optimal scalarizer function is difficult, if not impossible to find. As a solution, we instead determine which spectra to measure by training an auxiliary network to predict Im2spec reconstruction error. The curiosity-driven approach involves sampling regions with high values of the predicted error, so as to rapidly reduce the error of these models.
The paper presents two workflows: The first consists of an ensemble of Im2spec models that is used for spectral prediction, combined with an error model that trains on the spectral mismatch error. In the second method, the error model utilizes the latent space embeddings of an autoencoder to correlate with spectral mismatch. These algorithms, inspired by curiosity-driven reinforcement learning, actively sample spectra for which the structure-property relations have not yet been learned. We first demonstrate and optimize the efficacy of our methods on a pre-acquired dataset. Finally, we implement an algorithm on an atomic force microscope (AFM) to actively learn structure-property relationships in a ferroelectric thin film and discuss possible extensions.
Figure 1: Description of the spectral search method on pre-acquired dataset. (a) Spatial dependence of the spectral property that is correlated with the sample region shown on the left. The patches on the sample image serve as the structural inputs that is correlated to the spectral output. (b) An ensemble of different Im2spec models are trained on an initial set of data. (c) The best model is then used to predict the spectral output corresponding to the training image inputs. (d) Error model consists of the encoder and the latent embedding part of the Im2spec model. This is conjoined with a decoder that is used to train with the spectral mismatch (L1)error. The parameters of the encoder (and the latent embedding) are frozen while training the error model. (e) The error model is used to estimate errors for the image patches across the sample region. Subsequent sampling points are decided using the acquisition function and incorporated into the training set.
@@ III Results and Discussions
The structure-property relations described in this work correlate to the ferroelectric response of PbTiO3 samples measured using band excitation piezoresponse microscopy (BE-PFM). We use the structure information from the morphology data acquired from the AFM scanned images, while the property is measured using the spectra collected using band excitation piezoresponse spectroscopy (BEPS) data. This dataset is very similar to one captured and published earlier, and details about the measurement can be found elsewhereslautin2024bayesian .
Fig 1 illustrates the active learning workflow described in this section. Fig 1(a) shows sample dataset which shows spatial dependence of the local structure and its influence on the observed spectrum. Here, the local structure, indicated by the square patch, influences the spectrum measured in that region. We initially start by considering a training set where the inputs are the image patches (each patch of size (16 x 16) pixels) while the outputs are the spectra (256 points) corresponding to each patch. In principle, the choice of the patch size is a physics-based quantity which determines the extent of the local structure that affects the measured spectrum. In ferroelelectric measurements, this depends on the electrostatic and elastic fields. We also observe that the window size affects the correlative strength of the input and the output. This can be estimated by comparing the training and the validation loss for the different patch-sizes, results of which are shown in Fig. S1. Low values of the window size results in sub-optimal training while large window size can interfere with efficient learning and result in overfitting. Our choice of the patch window size (16 pixels) is in the optimal range, with low values of the validation loss.
In this workflow, we use an ensemble of Im2spec models to offer flexibility for variations in the training data (schematic as shown in Fig 1(b)). Each im2spec model consists of an encoder, a latent embedding layer, followed by a fully connected decoder. While the models are primarily based on the convolutional networks, variations in the architecture have been introduced to enable wider adaptability. A brief description of the encoder architectures used in the model set is provided in Table 1. In our workflow, we designate the size of the latent dimension as three. An initial dataset of the image patches is used to train the Im2spec models. During the training process, we implemented stochastic weighted average for stabilizing the model weights and for generalized spectral prediction. Once trained, the ”best model” of the ensemble is chosen based on the minimum validation loss, estimated over last 50 training epochs.
@@
The selected model is then used to predict the spectral output on the image inputs that were previously used for training (as shown in Fig 1(c)). This prediction is compared with the original spectrum, and the mismatch error is assigned to every image within the training set. We use the L1 error to quantify the spectral mismatch in this method. Fig 1(d) shows the error model where the Im2spec-encoder (which includes the latent embedding layer) is conjoined with a different set of decoder layers. During the error model training, the encoder part of the model is frozen while the decoder weights are updated. The next step involves the error prediction for the entire set of image patches across the sample region, as shown in Fig 1(e). The error predictions are used to compute the acquisition function to determine and sample the next set of spectral points in an iterative active learning fashion.
Our studies show that the best Im2spec model does not change frequently with minor changes in the training data set. Our code enables probabilistic triggering of ensemble training at selected iterations, helping to avoid redundant training steps. In the results described in this section, we perform ensemble training randomly over 10 % of the iterations (and the starting iteration). The remaining iterations involve model training using the pre-determined best-Im2spec-model.
Once we predict the errors for all set of the input patches, we use an acquisition function to sample the next data point. The acquisition function used in this method is an empirical equation and is given as:
@@
where $L_{j}$ is the L1 error normalized in the range [0, 1]. The $\beta$ parameter controls the rate of exploration and exploitation, while the prefactor $\lambda$ controls the smoothness of the acquisition function (higher $\lambda$ indicates better smoothness). The acquisition function varies monotonically with the error values for $\beta$ = 1 and inversely for $\beta$ = 0. This allows us to tune sampling from exploitation to exploration as we increase the $\beta$ hyperparameter from 0 to 1. Therefore the model with $\beta=1$ is exploration dominant, and is designated as curiosity driven model. In the results described in this section, we study model performance at the extreme, for $\beta$ = 0 and 1 while maintaining $\lambda$ = 0.1.
@@
The workflow starts with an initial dataset consisting of 245 image-spectrum pairs (20 % of the total dataset). Each iteration consists of two model training events - the im2spec ensemble models and the error model. As shown in Fig 1(e), we obtain the prediction of the error values at the end of each iteration. We use the acquisition to sample the next point in every iteration (an alternate method is batch sampling using the acquisition function). In the results described in this section, we study the model behavior over three hundred iterations of active learning.
Fig 2 shows the workflow results where we test the model for the $\beta$ parameter at 0 and 1. We compare the results of the model with a baseline model that trains on acquisitions based on random sampling. Fig 2(a) shows the error statistics for the three models. While the conservative model corresponding to $\beta$ = 0 shows low values of the spectral mismatch error, the curiosity model ($\beta=1$) shows acquisitions with higher values of the spectral mismatch error. In the curiosity model, at every iteration, new unfamiliar samples improves training over a diverse dataset leading to faster learning. To ascertain this behavior, we estimated the spectral mismatch error over the test set. The results described in Fig 2(b) show a steep reduction of the errors for the curiosity driven model.
Figures Fig 2(c), 2(d), and 2(e) show the acquisition points on the sample region for the random model, $\beta=0$, and $\beta=1$, respectively. We observe that the exploration for the $\beta=0$ is limited to the domains, and the $\beta=1$ acquires spectrum in the region of the domain walls and the defective regions of the sample, where the structure to spectral correlations are complex.
Figure 2: Results of the error prediction model in active learning on the pre-acquired BEPS data. Panel (a) shows the histogram of the error statistics using the acquisition function with random acquisition model, $\beta$ =0, and $\beta$ =1. The model with $\beta$ = 1 corresponds to the curiosity driven model that shows sampling of data associated with higher spectral mismatch errors. (b) Variation of spectral mismatch (L1) errors on the test-set for different models. The curiosity based model ($\beta$ = 1) shows steep reduction of the error across the iterations. (c), (d), and (e) shows the scatter points where the spectral measurements were acquired across the sample region for the random acquisition model, active learning with $\beta$ = 0, and $\beta$ = 1, respectively . All scale bars indicate a length of 100 nm.
@@
While we see extreme examples of exploitation and exploration for $\beta$ = 0 and 1, respectively, intermediate values of the $\beta$ can be used to balance exploration and exploitation. Fig S2 shows the results of the model and the acquisitions for $\beta=0.5$. This samples regions that correspond to both higher and lower values of the error prediction. The explorative performance of the model is therefore intermediate as shown the reduction of the test-set errors in Fig S2(b).
In the above analysis, and in the rest of the paper, we compare the performance of the model with the commonly used baseline i.e., random sampling. We believe this is reasonable baseline, especially while sampling from a multidimensional dataset. In section 3 of the supplementary material, we show a comparison different sampling techniques and its performance with respect to curiosity based active learning. We observe that the random sampling performs similar to other multidimensional sampling techniques. Nevertheless, curiosity-based active learning outperforms the other sampling techniques.
@@
In an encoder-decoder model, the latent representations that bridge the encoder and the decoder parts of the model determine the efficiency of the reconstruction. We study the latent embeddings to gain insights into the workings of the error model and to interpret the essential features that determine the model output. The latent distributions of the model predictions are described in Fig 3 for the active learning process. Fig 3(a)-(c) represents the latent space distributions for the random model. Fig 3(a) is the latent distribution with the red scatter points, which denote explorations during the active learning process, sampled uniformly. Fig 3(b) shows the latent space clustered into 3 classes, and the corresponding correlation to the real space is shown in Fig 3(c). It is to be noted that the acquisition strategy influences the evolution of the training set, the model weights, and therefore the latent representations. In the random sampling we see uniform sampling across the clusters. Fig 3(d)-(f) shows the latent distribution for $\beta=0$. The conservative nature model is reflected in limited exploration that are localized at the high-density region of the latent space. A similar analysis is performed for exploration related to $\beta$ = 1, shown in Fig 3(g)-(i). Here, higher exploration has resulted in a dispersed latent distribution. Further exploration points are comparatively sampled in the sparse region of the latent space. In the real space mapping (Fig 3(i)), this translates to acquisition in the complementary areas (when compared to $\beta=0$) and corroborates with the data shown in Fig 2(e).
@@
Figure 3: Latent space representations for the active learning-based sampling. Panel (a)-(c) shows the results of the model for randomly sampled model (a) Latent distribution of the model prediction for all image-patch inputs. The red scatter points are the explored samples during the active learning process. (b) depicts the latent distribution clustered into three components using k-means clustering. (c) Correlation to the real-space coordinates corresponding each of the points in the latent space. Panel (d)-(f) shows the similar results for active learning-based sampling with $\beta$ = 0. Figures (g)-(i) show latent distribution results for active learning-based sampling for $\beta$ = 1. While $\beta=0$ shows predominant sampling in the high dense regions, $\beta=1$ shows sampling in the dispersed regions of the latent space.
@@
The results of this section describe the error prediction methods in conjunction with the acquisition function, where the $\beta$ parameter is used to control the degree of exploration/exploitation. At higher $\beta=1$, the model is curiosity driven and actively seeks unfamiliar samples in the spatial regions of higher predicted error. This allows for diversity within the training set to better learn structure-spectral correlations.
@@
The latent embedding show uniquely different distributions based on the model and the acquisition strategy. These embeddings serve as compressed, structured representations of input data, capturing essential features of the input images. Given this knowledge, in the subsequent section, we implement a generalized methodology to extract latent representations from an autoencoder while efficiently sampling points from the latent space for active learning based acquisitions.
@@ III.2 Autoencoder-based error model
Figure 4: Diagram of Curiosity Algorithm implementation with Im2spec used in conjunction with the Autoencoder-based error model
This section describes the autoencoder based error model. Here we use a similar experimental dataset - PFM-based experimental data on a 200 nm (110) $PbTiO_{3}$ thin film sample grown on $SrTiO_{3}$. The structural information is contained in the image patches (patch size (11$\times$11 pixels)) with switching spectroscopy spectra (spectrum length: 64) captured at low frequency (off-resonance) using an interferometric atomic force microscope from Oxford Instruments (Vero). The baseline curiosity algorithm works as follows: After the sample is imaged an autoencoder is trained on all image patches. Then spectra are acquired on a small number of random initialization points, which are used to train the Im2spec model. The error predictor is then trained on the image patch latent encodings and the Im2spec mean squared error (MSE) for the initial points. Then, a forward pass through the error predictor is performed for all image patch encodings. The spectra of the point with the highest predicted error is then sampled. This continues iteratively where Im2spec and the error predictor are trained on updated training dataset. The overall workflow is illustrated in Fig 4. It should be noted that the error predictor model utilizes dropout to provide an estimate of the uncertainty on the prediction.
@@
Figure 6: Minimum loss achieved by Im2spec with Curiosity Algorithm vs Random sampling, for 30 trials. The difference between the means is statistically significant
Figure 7: MSE for ten highest error points achieved by Im2spec with MC Dropout Curiosity Algorithm vs Random sampling. The difference between the means is statistically significant.
Figure 8: Trial of Curiosity Algorithm real-time on PFM microscope for Spec2im and Im2spec. (a) Exploration path of Spec2im Curiosity Algorithm, (b) predicted Spec2im error after final measurement iteration, (c) a polarization ground-truth image-patch, and (d) corresponding Spec2im prediction. (e) Exploration path of Im2spec Curiosity Algorithm, (f) predicted Im2spec error after final measurement iteration, (g) a ground-truth hysteresis loop and corresponding Im2spec prediction. Scale bar in the images indicate a length of 100 nm.
@@
In addition, we tested a modified curiosity algorithm which, in addition to latent space exploration reward, samples based on Im2spec Monte-Carlo Dropout (MCD) uncertainty during the exploration phase. While the addition of MCD uncertainty did not directly improve Im2spec loss (the Im2Spec loss was overall not statistically different from a random sampling strategy, in this case), it did reduce the Im2spec MSE for the ten highest error points (Fig. 7). This behavior suggests that enhancing exploration with MCD helps train Im2spec on points with poorly understood structure property relationships, but are not abundantly represented in the sample data, as opposed to points with low error, but are highly represented in the sample data. This, however, has the downside of slower convergence on the whole dataset compared to the original (non-MCD) case. It should be noted that one of the challenges of this algorithm is that there may exist points that continue to contain high errors regardless of the number of training data points, if there are minimal structure-property correlations in these points (for example, if there is only noise in these areas). For such instances, the algorithm should be modified to avoid trapping in these learning plateaus, and strategies can include either direct human intervention, injected noise in the action space, or simple methods such as avoiding similar image patches to past samples if the loss is not decreasing beyond a simple threshold.